QUESTION9 2 points Save Answer Determine how many different committees can be formed if 3 members must be teachers and 2 members must be students. TTI Path: P words:0 QUESTION 10 2points Save Answer A class committee is to be made that consists of 2 freshmen, 2 sophomores, 3 juniors, and 3 seniors. If students are being randomly chosen from groups of 12 freshmen, 17 sophomores, 16 juniors, and 20 seniors, how many possible committees can be formed? Words QUESTION 11 2 points Save Answer The probability of rain on any given day is 3. What is the probability of at most one day of rain during the next three days? Use the equation editorx to enter your answer as a fraction
Answer #1

Q9- Incomplete.

Q10-12C2(selecting 2 freshmen from 12)

17c2- 2 sophomores

16c3- 3 juniors

20c3- 3 seniors

so Toal would be 12c2.17c2.16c3.20c3.

Q11- probability of no rain =1/3

so on atmost 1 day rain

3*1/3*1/3*2/3+1/3*1/3*1/3=7/27